### v3

# Jouko Väänänen Helsinki and Amsterdam Ur Ya’ar

Abstract
If $\mathcal{L}$ is an abstract logic (a.k.a. model theoretic logic), we can define the inner model $C(\mathcal{L})$ by replacing first order logic with $\mathcal{L}$ in Gödel’s definition of the inner model $L$ of constructible sets. Set theoretic properties of such inner models $C(\mathcal{L})$ have been investigated recently and a spectrum of new inner models is emerging between $L$ and $\operatorname{HOD}$. The topic of this paper is the effect on $C(\mathcal{L})$ of a slight modification of $\mathcal{L}$ i.e. how sensitive is $C(\mathcal{L})$ on the exact definition of $\mathcal{L}$? The $\Delta$-extension $\Delta(\mathcal{L})$ of a logic is generally considered a “mild” extension of $\mathcal{L}$. We give examples of logics $\mathcal{L}$ for which the inner model $C(\mathcal{L})$ is consistently strictly smaller than the inner model $C(\Delta(\mathcal{L}))$, and in one case we show this follows from the existence of $0^{\sharp}$.
1 Introduction
This paper is part of a larger project focusing on applications of strong logics in set theory. A pioneering result in this direction is Menachem Magidor’s characterization of Löwenheim-Skolem type properties as well as compactness properties of second order logic in terms of supercompact and extendible cardinals [Mag71]. Another example is the so-called Magidor-Malitz quantifier ([MM77]). A more recent example is the project of “extended constructibility" ([KMV21, KMV25]).
In [KMV21] an inner model of set theory $C(\mathcal{L})$ was associated with any abstract logic $\mathcal{L}$ (all relevant notions are defined below). For many logics $\mathcal{L}$ this inner model is just Gödel’s inner model $L$ of constructible sets. But consistently, and also assuming large cardinals, $C(\mathcal{L})$ can be very different from Gödel’s $L$ [KMV21]. For example, if $\mathcal{L}$ is the stationary logic of [BKM78] and we assume that there is a proper class of Woodin cardinals, the resulting inner model $C(\mathcal{L})$, denoted $C(\mathtt{aa})$, has a proper class of measurable cardinals [KMV25]. Associated with any abstract logic $\mathcal{L}$ there is an extension $\Delta(\mathcal{L})$ of $\mathcal{L}$, introduced in [Bar74], see also [MSS76]. We give the definition below. It is generally considered that the difference between $\Delta(\mathcal{L})$ and $\mathcal{L}$ is minor, from the model theoretic point of view, in the sense that $\Delta$ preserves a lot of model theoretic properties of the logic $\mathcal{L}$. In this paper we compare the inner models $C(\mathcal{L})$ and $C(\Delta(\mathcal{L}))$ and show that these can be different (of course, if $V=L$ they are the same, so we look for consistency results and results under large cardinal assumptions). This shows that the set-theoretical power of $\Delta(\mathcal{L})$ and $\mathcal{L}$ can be quite different even if they greatly resemble each other from a model theoretic view point. We begin by recalling definitions and facts we will use throughout the paper.
1.1 Abstract Logics and the $\Delta$-operation
We present the basic notions we require from the theory of abstract logics. For further information, see [BF85].
An abstract logic $\mathcal{L}$ consists of a pair $(S,T)$ where, given a vocabulary $\tau$, $S$ defines the class of sentences of $\mathcal{L}$ in the vocabulary $\tau$, and $T$ defines the satisfaction relation between $\tau$-structures and members of $S$. We usually say simply “a logic” instead of “abstract logic”, abuse notation by writing e.g. $\varphi\in\mathcal{L}$ instead of $S(\varphi)$, and write $\mathcal{M}\models_{\mathcal{L}}\varphi$ instead of $T(\mathcal{M},\varphi)$. We may consider formulas with free variables by adding constant symbols to the vocabulary. While this definition is quite general, prominent examples of abstract logics are concretely given by supplementing the syntax of first-order logic with additional quantifiers and extending the semantics with appropriate clauses for these quantifiers (some examples are presented below). A core idea of the field of abstract logics, giving it the synonym “Model Theoretic” logics, is looking at classes of models as defining logics. We recall some definitions.
Definition 1.
A model class is any class of models of a fixed vocabulary $\tau$ which is closed under isomorphisms. Every sentence $\phi$ of a logic $\mathcal{L}$ defines a model class, namely the class of models satisfying $\phi$. Suppose $\mathcal{K}$ is a model class of vocabulary $\tau$. $\mathcal{K}$ is said to be:
  • $\mathcal{L}$-definable if there is a sentence $\phi$ of $\mathcal{L}$ such that $\mathcal{K}$ is the class of models satisfying $\phi$;
  • $\Sigma(\mathcal{L})$-definable if there is a bigger vocabulary $\tau^{\prime}$, possibly many-sorted, and a sentence $\phi$ of $\mathcal{L}$ in the vocabulary $\tau^{\prime}$ such that $\mathcal{K}$ is the class of reducts of models of $\phi$ to $\tau$;
  • $\Pi(\mathcal{L})$-definable if its complement (in the class of all models of vocabulary $\tau$) is $\Sigma(\mathcal{L})$-definable;
  • $\Delta(\mathcal{L})$-definable if both $\mathcal{K}$ and its complement are $\Sigma(\mathcal{L})$-definable (in other words, if it is both $\Sigma(\mathcal{L})$- and $\Pi(\mathcal{L})$-definable).
A logic $\mathcal{L}^{\prime}$ extends another logic $\mathcal{L}$, in symbols $\mathcal{L}\leq\mathcal{L}^{\prime}$, if every definable model class of $\mathcal{L}$ is also a definable model class of $\mathcal{L}^{\prime}$. The logics $\mathcal{L}$ and $\mathcal{L}^{\prime}$ are equivalent, in symbols $\mathcal{L}\equiv\mathcal{L}^{\prime}$, if $\mathcal{L}\leq\mathcal{L}^{\prime}$ and $\mathcal{L}^{\prime}\leq\mathcal{L}$.
The $\Delta$-extension $\Delta(\mathcal{L})$ of an abstract logic $\mathcal{L}$ is the $\leq$-least logic $\mathcal{L}^{*}$ such that every $\Delta(\mathcal{L})$-definable model class is $\mathcal{L}^{*}$ definable.
In the most abstract setting, $\Delta(\mathcal{L})$ can be defined as the logic whose sentences are simply the $\Delta(\mathcal{L})$-definable model classes, so the satisfaction relation is just the membership relation. More concretely, we can follow [BF85, Definition II.7.2.3] and let the sentences of $\Delta(\mathcal{L})$ in a vocabulary $\tau$ be pairs $(\phi,\psi)$ such that $\phi$ and $\psi$ are $\mathcal{L}$-sentences in vocabularies extending $\tau$, and:
[…]
Intuitively speaking, $\Delta(\mathcal{L})$-definability is a form of implicit definability, and is related to the notion of interpolation:
Definition 2.
A logic $\mathcal{L}$ is said to have the Craig Interpolation Property if every two disjoint $\Sigma(\mathcal{L})$-definable model classes (of the same vocabulary) can be separated by a definable model class of $\mathcal{L}$. $\mathcal{L}$ has the Suslin-Kleene Interpolation Property if every two complementary $\Sigma(\mathcal{L})$-definable classes can be separated by a definable class of $\mathcal{L}$.
In other words, $\mathcal{L}$ has the Suslin-Kleene Interpolation Property iff every $\Delta(\mathcal{L})$-definable class is actually definable iff $\Delta(\mathcal{L})=\mathcal{L}$. Clearly every logic with the Craig Interpolation Property has the Suslin-Kleene Interpolation Property, but the converse is generally false ([Hut76, Theorem 3.1]).
$\Delta(\mathcal{L})$ can be considered a minor extension of $\mathcal{L}$ in the sense that it preserves many model theoretic properties of $\mathcal{L}$:
Fact 3 ([MSS76]).
Let $\mathcal{L}$ be an abstract logic. Then
  • $\Delta(\mathcal{L})$ is $(\kappa,\lambda)$-compact (for infinite cardinals $\kappa\geq\lambda$) if $\mathcal{L}$ is;
    Definition 4.
    A logic $\mathcal{L}$ is said to be $(\kappa,\lambda)$-compact whenever given a set of $\mathcal{L}$-sentences $\Sigma$ of cardinality $\kappa$, if each subset of $\Sigma$ of cardinality $\lambda$ has a model, then $\Sigma$ has a model. A logic is $\kappa$-compact if it is $(\kappa,\aleph_{0})$-compact.
  • $\Delta(\mathcal{L})$ has the same Löwenheim-number as $\mathcal{L}$;
    Definition 5.
    The Löwenheim-number of $\mathcal{L}$ is the smallest cardinal $\kappa$ such that if an arbitrary sentence of $\mathcal{L}$ has any model, the sentence has a model of cardinality no larger than $\kappa$.
  • $\Delta(\mathcal{L})$ is axiomatizable if $\mathcal{L}$ is;
  • $\Delta(\mathcal{L})$ is the minimal extension of $\mathcal{L}$ to a logic with the Suslin-Kleene Interpolation Property;
  • $\Delta(\Delta(\mathcal{L}))=\Delta(\mathcal{L})$.
A weakness of the $\Delta$-operation is that its syntax depends, a priori, on set theory. Depending on the logic $\mathcal{L}$, the question whether a pair $(\phi,\psi)$ is a sentence of $\Delta(\mathcal{L})$, i.e. whether the equivalence (1) above holds, may be non-absolute in set theory and of high set theoretical complexity. This is in sharp contrast to the absoluteness of the syntax of most logics that one encounters, e.g. first order logic and its extensions by generalized quantifiers. In fact, we exploit a non-absolute case of (1) in our proof of Theorem 16 below. One consequence of the potential non-absoluteness of the syntax of $\Delta(\mathcal{L})$ is that the usual method of proving that the Axiom of Choice (AC) holds in models of the form $C(\mathcal{L})$ (Fact 12) cannot be used, as it requires having a recursive set of formulas. However our results do not require AC to hold in the models we discuss, so we leave the question whether or not it holds open.
1.2 Inner models from extended logics
We review some of the definitions and facts from [KMV21] regarding constructing inner models using extended logics.
Definition 6.
Let $\mathcal{L}$ be a logic (extending first-order logic).
  1. 1.
    Let $M$ be a set. Define:
    $\operatorname{Def}_{\mathcal{L}}\left(M\right)=\left\{\left\{a\in M\mid\left(M,\in\right)\models_{\mathcal{L}}\varphi(a,\bar{b})\right\}\mid\varphi\in\mathcal{L};\,\bar{b}\in M^{<\omega}\right\}$
  2. 2.
    $C\left(\mathcal{L}\right)$, the class of $\mathcal{L}$-constructible sets, is defined by induction:
    $L^{\prime}_{0}$ $=\varnothing$
    $L^{\prime}_{\alpha+1}$ $=\operatorname{Def}_{\mathcal{L}}\left(L^{\prime}_{\alpha}\right)$
    $L_{\beta}^{\prime}$ $=\bigcup_{\alpha<\beta}L_{\alpha}^{\prime}\,\,\,\text{for limit $\beta$}$
    $C\left(\mathcal{L}\right)$ $=\bigcup_{\alpha\in Ord}L^{\prime}_{\alpha}$
Note that if $\mathcal{L}$ is first-order logic, then we get exactly Gödel’s $L$. It is also easy to see that if $\mathcal{L}\leq\mathcal{L}^{\prime}$, then $C(\mathcal{L})\subseteq C(\mathcal{L}^{\prime})$, but, of course, not conversely.
Fact 7 ([KMV21] Proposition 2.3).
For any logic $\mathcal{L}$, $C\left(\mathcal{L}\right)$ is a transitive model of $\mathrm{ZF}$ containing all the ordinals.
Fact 8 ([KMV21] Proposition 2.9).
If the logic $\mathcal{L}$ is defined by hereditarily ordinal definable parameters, and each of its formulas is hereditarily ordinal definable (with finitely many free variables), then $C(\mathcal{L})\subseteq\operatorname{HOD}$.
Definition 9.
Suppose $A$ is any class and $T$ is any theory in the language of set theory. A logic $\mathcal{L}^{*}$ is $T$-absolute if there are a $\Sigma_{1}$-predicate $S_{1}(x)$, a $\Sigma_{1}$-predicate $T_{1}(x,y)$, and a $\Pi_{1}$-predicate $T_{1}^{\prime}(x,y)$ such that $\varphi\in\mathcal{L}^{*}\Leftrightarrow S_{1}(\varphi)$, $M\models\varphi\Leftrightarrow T_{1}(M,\varphi)$ and $T\vdash\forall x\forall y\left(S_{1}(x)\rightarrow\left(T_{1}(x,y)\leftrightarrow T_{1}^{\prime}(x,y)\right)\right)$. If parameters from a class $A$ are allowed, we say that $\mathcal{L}^{*}$ is absolute with parameters from $A$.
Furthermore, $\mathcal{L}^{*}$ has $T$-absolute syntax if its sentences are (coded as) natural numbers and there is a $\Pi_{1}$-predicate $S_{1}^{\prime}(x)$ such that $T\vdash\forall x\left(S_{1}(x)\leftrightarrow S_{1}^{\prime}(x)\right)$. We may allow parameters as above.
Fact 10 ([KMV21] Theorem 3.4).
Suppose $\mathcal{L}^{*}$ is $\mathrm{ZFC}$-absolute with parameters from $L$, and the syntax of $\mathcal{L}^{*}$ is $\mathrm{ZFC}$-absolute with parameters from $L$. Then $C\left(\mathcal{L}^{*}\right)=L$.
Recall that an admissible set is a transitive set $X$ such that $\left<X,\in\right$> satisfies $\mathrm{KP}$ – the Kripke-Platek axioms of set theory.
Definition 11.
A logic $\mathcal{L}^{*}$ is adequate to truth in itself if for all finite vocabularies $\tau$ there is function $\varphi\mapsto\ulcorner\varphi\urcorner$ from all formulas $\varphi(x_{1},\ldots,x_{n})\in\mathcal{L}^{*}$ in the vocabulary $\tau$ into $\omega$, and a formula $\operatorname{Sat}_{\mathcal{L}^{*}}(x,y,z)$ in $\mathcal{L}^{*}$ such that:
  1. 1.
    The function $\varphi\mapsto\ulcorner\varphi\urcorner$ is one-to-one and has a recursive range.
  2. 2.
    For all admissible sets $M$, formulas $\varphi$ of $\mathcal{L}^{*}$ in the vocabulary $\tau$, structures $\mathcal{N}\in M$ in the vocabulary $\tau$, and $a_{1},\ldots,a_{n}\in N$, the following conditions are equivalent:
    1. (a)
      $\langle M,\in\rangle\models\operatorname{Sat}_{\mathcal{L}^{*}}\left(\mathcal{N},\ulcorner\varphi\urcorner,\left\langle a_{1},\ldots,a_{n}\right\rangle\right)$
    2. (b)
      $\mathcal{N}\models\varphi\left(a_{1},\ldots,a_{n}\right)$.
We may admit ordinal parameters in this definition.
Fact 12 ([KMV21] Proposition 2.7).
If $\mathcal{L}^{*}$ is adequate to truth in itself, then $C\left(\mathcal{L}^{*}\right)$ satisfies the Axiom of Choice.
2 Initial results
The standard example of the use of the $\Delta$-operation is the result of [Bar74] to the effect that $\Delta(\mathcal{L}(Q_{0}))=\mathcal{L}_{\mbox{\tiny HYP}}$, where $Q_{0}$ is the quantifier asserting “there are $\geq\aleph_{0}$ many…” and $\mathcal{L}_{\mbox{\tiny HYP}}$ is the smallest admissible fragment of $\mathcal{L}_{\omega_{1}\omega}$. This result implies that $\mathcal{L}(Q_{0})\neq\mathcal{L}_{\mbox{\tiny HYP}}$ (as Barwise [Bar74] shows in the proof of Corollary 4.3). By fact 10 $C(\mathcal{L}(Q_{0}))=C(\mathcal{L}_{\mbox{\tiny HYP}})=L$, as both $\mathcal{L}(Q_{0})$ and $\mathcal{L}_{\mbox{\tiny HYP}}$ are absolute logics (by [Bar72]). In conclusion, $\Delta(\mathcal{L}(Q_{0}))\neq\mathcal{L}(Q_{0})$ but $C(\Delta(\mathcal{L}(Q_{0})))=C(\mathcal{L}(Q_{0}))$.
Another example of this phenomenon is second order logic $\mathcal{L}^{2}$. Assuming the Axiom of Choice, the inner model $C(\mathcal{L}^{2})$ is equal to $\operatorname{HOD}$ ([MS71]), and therefore also $C(\Delta(\mathcal{L}^{2}))$ is equal to $\operatorname{HOD}$ by Fact 8 (note that the definition of $\Delta(\mathcal{L}^{2})$ doesn’t require any parameters which aren’t hereditarily ordinal definable). Hence $C(\Delta(\mathcal{L}^{2}))=C(\mathcal{L}^{2})$, but $\Delta(\mathcal{L}^{2})\neq\mathcal{L}^{2}$ by [Cra] (which essentially shows that the satisfaction relation for $\mathcal{L}^{2}$ is a counterexample). Note that $C(\mathcal{L}^{2})$ may be a proper subclass of $\operatorname{HOD}$ in the absence of AC ([Szc77]).
Thus we have established:
Proposition 13.
There are logics $\mathcal{L}$ such that $\Delta(\mathcal{L})\neq\mathcal{L}$ but provably, $C(\Delta(\mathcal{L}))=C(\mathcal{L})$.
We now show that there are logics where the $\Delta$-operation might not preserve the inner model $C(\mathcal{L})$. Let $H$ be the Henkin quantifier [Hen61]
(2) $Hxyuv\ \phi(x,y,u,v,\vec{w})=\left(\begin{array}[]{cc}\forall x&\exists y\\ \forall u&\exists v\end{array}\right)\phi(x,y,u,v,\vec{w})$
the meaning of which is
(3) $\exists f\exists g\forall x\forall u\,\phi(x,f(x),u,g(u),\vec{w}).$
Let us use $C(H)$ to denote the inner model $C(\mathcal{L}(H))$. Recall that $\operatorname{HOD}_{1}$ is defined as $C(\Sigma^{1}_{1})$ ([KMV21]), where $\Sigma^{1}_{1}$ is existential second order logic, i.e. the logic of formulas of second order logic which start with an existential second order quantifier and then continue with a first order formula.
Theorem 14.
$C(H)=\operatorname{HOD}_{1}$.
Proof.
By [Wal70, Corollary 4.6], in the presence of a pairing function any $\Sigma^{1}_{1}$-formula can be expressed in terms of the Henkin quantifier $H$. This implies $\operatorname{HOD}_{1}\subseteq C(H)$. Conversely, any formula starting with $H$ can be translated via (3) into a $\Sigma^{1}_{1}$-form. This makes it possible to prove by induction on the structure of the formula defining a set $X\in C(H)$ that $C(H)\subseteq\operatorname{HOD}_{1}$. For example if $X$ is defined using $\neg Hxyuv\ \phi$, then its complement relative to some level of the construction is, by induction, in $\operatorname{HOD}_{1}$, and so $X$ will enter $\operatorname{HOD}_{1}$ in the next level. ∎
Since our $\Delta$-operation is many-sorted, $\Delta(\mathcal{L}(H))$ contains full second order logic [KL79]. The logic $\mathcal{L}(H)$ is not $\Delta$-closed, i.e. $\Delta(\mathcal{L}(H))\neq\mathcal{L}(H)$ ([GH76]), but we do not need this here. By [KMV21, Theorem 7.9] it is consistent, relative to the consistency of $\mathrm{ZF}$, that $\operatorname{HOD}_{1}\neq\operatorname{HOD}$. Hence:
Theorem 15.
$Con(\mathrm{ZF})$ implies $Con(C(\mathcal{L}(H))\neq C(\Delta(\mathcal{L}(H))))$.
3 The quantifier $Q_{1}$
As a second example, let $Q_{1}$ be the quantifier asserting “there are $\geq\aleph_{1}$ many…”. It is well-known that $\Delta(\mathcal{L}(Q_{1}))\neq\mathcal{L}(Q_{1})$, as demonstrated by the property of an equivalence relation having uncountably many uncountable classes (see [BF85, Chapter II section 7.2]). This raises the question, what is the relationship between $C(\mathcal{L}(Q_{1}))$ and $C(\Delta(\mathcal{L}(Q_{1})))$?
Theorem 16.
$Con(\mathrm{ZF})$ implies $Con(C(\mathcal{L}(Q_{1}))\neq C(\Delta(\mathcal{L}(Q_{1}))))$.
The proof will use the following machinery. Recall that an Aronszajn tree is a tree of height $\omega_{1}$ such that all its levels are countable, and it has no uncountable chain (a branch); a Suslin tree is an Aronszajn tree which has no uncountable antichain; and an Aronszajn tree $S$ is called special if there is a specializing function – a function $c:S\to A$ such that $A$ is countable, and for every $q,p\in S$, if $c(q)=c(p)$ then $q,p$ are incompatible in $S$. Note that a special Aronszajn tree must have an uncountable chain, so cannot be Suslin. In general, there might be Aronszajn trees that are neither special nor Suslin.
A Suslin tree can be considered as a forcing notion which has the countable chain condition (ccc), and adds a generic uncountable chain to itself. Furthermore, if $S,T$ are Suslin trees such that the tree product $S\otimes T$ is Suslin as well, then forcing with one preserves the Suslinity of the other. Note that, since the tree product $S\otimes T$ is dense in the forcing product $S\times T$, this corresponds to the following well-known fact (see [Jec07, Lemma 16.6]):
Fact 17.
[…]
To summarize, we have the following properties:
Lemma 18.
If $T,S$ are Suslin trees such that $T\otimes S$ is Suslin, then
  1. 1.
[…]
  2. 2.
[…]
Suslin trees satisfying the conclusion of the lemma are called independent. By [AS93, Theorem 7.3], there is a sequence of Suslin trees $\mathcal{T}=\left\langle T_{n}\mid n<\omega\right\rangle\in L$ such that for every distinct $i_{1},...,i_{k}\in\omega$, the product tree $T_{i_{1}}\otimes...\otimes T_{i_{k}}$ is Suslin. To summarize, for every $i\in\omega$:
[…]
  • $T_{i},\mathcal{S}(T_{i})\Vdash$ “for every distinct $i_{1},...,i_{k}\in\omega\smallsetminus\{i\}$, the product tree $T_{i_{1}}\otimes...\otimes T_{i_{k}}$ is Suslin”.
So, all in all, we can either force a branch or specialize $T_{i}$ without affecting the Suslinity of any of the other $T_{j}$s or the independence between them. Note that all posets used, as well as their finite products, are ccc, and as a consequence a finite support product of them preserves $\omega_{1}$.
Proof of theorem 16.
By [KMV21, section 3], $\mathrm{ZF}\vdash C(\mathcal{L}(Q_{1}))=L$. So we wish to construct a model in which $C(\Delta(\mathcal{L}(Q_{1})))\neq L$.
Let $a\subset\omega$ be a Cohen real over $L$. Using a finite support product, we force over $L[a]$ to add a branch to any $T_{i}$ for $i\in a$ (using $T_{i}$ itself) and specialize any $T_{j}$ for $j\notin a$ (using $\mathcal{S}(T_{j})$). As we noted, this product preserves $\omega_{1}$. Then we force Martin’s Axiom ($\mathsf{MA}$) in the standard way (as in [ST71]) over the resulting model, and denote the final model $W$. Note that still, $\omega_{1}^{W}=\omega_{1}^{L}$. We now work in $W$. As a consequence of $\mathsf{MA}$, any Aronszajn tree in $W$ is special ([BMR70]), and we have
$a=\left\{i<\omega\mid T_{i}\text{ is not Aronszajn}\right\}.$
Claim 19.
$a\in C(\Delta(\mathcal{L}(Q_{1})))$.
Proof.
[…]
$\exists B\left(\forall x,y\in T(Bx\land By\to(x\leq y\lor y\leq x))\land Q_{1}xBx\right).$
Since by $\mathsf{MA}$ all Aronszajn trees are special, an $\aleph_{1}$-tree is not Aronszajn also iff it is not special. But being a special tree is $\Sigma(\mathcal{L}(Q_{1}))$:  [^1] Here we use $\omega$ as a parameter from $L^{\prime}_{\alpha}$, but this can be avoided as we have $Q_{1}$ at our disposal, and all we need is that the range of $F$ is countable.
$\exists F:T\to\omega\left(\forall x,y\in T(f(x)=f(y)\to x\bot y)\right).$
So under $\mathsf{MA}$, the property of being an Aronszajn tree is $\Delta(\mathcal{L}(Q_{1}))$. Hence indeed, in $W$, $a\in C(\Delta(\mathcal{L}(Q_{1})))$. ∎
To conclude, as $a$ is Cohen generic over $L$, we obtain
$W\models C(\mathcal{L}(Q_{1}))\neq C(\Delta(\mathcal{L}(Q_{1}))).\qed$
4 Symbiosis
As a next step it would be natural to ask whether certain properties of a logic $\mathcal{L}$ might imply that $C(\mathcal{L})=C(\Delta(\mathcal{L}))$ even though $\mathcal{L}\neq\Delta(\mathcal{L})$. One such candidate is the notion of symbiosis [Vää79]. The idea of symbiosis is to establish a close connection between an abstract logic and predicates of set theory. The motivation for this is the fact that many abstract logics seem to depend to a great extent on set theory, so it makes sense to ask exactly how strong is this dependence. A paramount example is second order logic and its symbiosis with the power-set operation of set theory.
Definition 20.
A set-theoretic relation $R$ and a logic $\mathcal{L}^{*}$ are symbiotic if the following conditions are satisfied:
  1. 1.
    Every $\mathcal{L}^{*}$-definable model class is $\Delta_{1}(R)$-definable.
  2. 2.
    Every $\Delta_{1}(R)$-definable model class is $\Delta(\mathcal{L}^{*})$-definable.
As noted in [BV16], the second clause can be replaced with
  1. $\text{2}^{*}$.
    The model class
    $Q_{R}:=\left\{\mathcal{A}\mid\mathcal{A}\cong\left(M,\in,\vec{a}\right),M\text{ transitive and }R(\vec{a})\right\}$
    is $\Delta(\mathcal{L}^{*})$-definable.
Proposition 21.
$\mathcal{L}(Q_{1})$ is not symbiotic with any set theoretic predicate.
Proof.
Symbiosis implies that the well-ordering quantifier $\mathrm{W}$, defined by
$\mathcal{M}\models\mathrm{W}xy\varphi(x,y)\iff\{(a,b)\in M^{2}\mid\mathcal{M}\models\varphi(a,b)\}\text{ is a well-order},$
is $\Delta(\mathcal{L})$-definable (cf. [Vää79, pg. 400]). But the $\Delta$-operation preserves compactness (Fact 3), and $\mathcal{L}(Q_{1})$ is $\aleph_{0}$-compact (see e.g. [BF85, II.3]), so $\Delta(\mathcal{L}(Q_{1}))$ is $\aleph_{0}$-compact as well, which in particular implies that the well-ordering quantifier is not definable. So $\mathcal{L}(Q_{0})$ cannot be symbiotic with any set-theoretic relation. ∎
However this can be easily amended by adding $\mathrm{W}$ to the logic, since $C(\mathcal{L}(Q_{1},\mathrm{W}))=L$ as well (again by [KMV21, section 3]), and we do have:
Proposition 22.
The logic $\mathcal{L}(Q_{1},\mathrm{W})$ is symbiotic with the predicate $R_{\aleph_{1}}$ (where $R_{\aleph_{1}}(a)$ $\iff$ $\left|a\right|\geq\aleph_{1}$).
Proof.
Denote $\mathcal{L}=\mathcal{L}(Q_{1},\mathrm{W})$ and $R=R_{\aleph_{1}}$. We verify clauses $1$ and $2^{*}$ above.
1) Every $\mathcal{L}$-definable model class is $\Delta_{1}(R)$-definable:
Let $\mathcal{K}=\left\{\mathcal{A}\mid\mathcal{A}\models_{\mathcal{L}}\varphi\right\}$ for some $\varphi\in\mathcal{L}$. Denote by $\mathrm{ZFC}_{n}^{-}$ the axioms of $\mathrm{ZFC}$ without the Power Set Axiom, and with the axiom schemata of Separation and Collection restricted to $\Sigma_{n}$ formulas. $\mathrm{ZFC}_{n}^{-}$ is finitely axiomatizable. Then there is some $n<\omega$ such that the following are equivalent:
  1. a)
    $\mathcal{A}\in\mathcal{K}$
  2. b)
    $\exists M$ $\big($ $M$ is transitive $\land$ $M\models\mathrm{ZFC}_{n}^{-}$ $\land$ $\forall a\in M$ ( $R(a)$ $\leftrightarrow$ $M\models\left|a\right|\geq\aleph_{1}$ ) $\land$ $\mathcal{A}\in M$ $\land$ $M\models$$\mathcal{A}\models_{\mathcal{L}}\varphi$$\big)$.
  3. c)
    $\forall M$ $\big[$$\big($ $M$ is transitive $\land$ $M\models\mathrm{ZFC}_{n}^{-}$ $\land$ $\forall a\in M$ ( $R(a)$ $\leftrightarrow$ $M\models\left|a\right|\geq\aleph_{1}$ ) $\land$ $\mathcal{A}\in M$ $\big)\to$ $M\models$$\mathcal{A}\models_{\mathcal{L}}\varphi$$\big]$ .
Note that since we consider transitive $M$’s, they are correct about the notion of well-order.
2) The model class $Q_{R}=\left\{\mathcal{A}\mid\mathcal{A}\cong(M,\in,a),M\text{ transitive and }|a|\geq\aleph_{1}\right\}$, in the language $\left(E,c\right)$ where $E$ is binary and $c$ is a constant, is $\Delta(\mathcal{L})$-definable:
  • Consider the class of (2-sorted) models $\mathcal{A}=\left(A,B,E^{\mathcal{A}},c^{\mathcal{A}},<^{\mathcal{A}},F^{\mathcal{A}}\right)$ such that $E^{\mathcal{A}}$ is a binary relation on $A$ satisfying the Extensionality Axiom, $c^{\mathcal{A}}\in A$, $^{\mathcal{A}}$ is a well-order on $B$ (definable using $\mathrm{W}$), $F^{\mathcal{A}}:\left(A,E^{\mathcal{A}}\right)\to(B,<^{\mathcal{A}})$ is an order-preserving function (a rank function) and $R(c^{\mathcal{A}})$ holds. It is elementary in $\mathcal{L}(Q_{1},\mathrm{W})$, and its projection yields the class of well-founded models $\left(A,E^{\mathcal{A}},c^{\mathcal{A}}\right)$ such that $R(c^{\mathcal{A}})$. The transitive collapse of such a model is of the form $(M,\in,\bar{c})$ where $\bar{c}$ is uncountable. So this is exactly $Q_{R}$.
  • Consider the class of (2-sorted) models $\mathcal{A}=\left(A,B,E^{\mathcal{A}},c^{\mathcal{A}},<^{\mathcal{A}},G^{\mathcal{A}}\right)$ such that $E^{\mathcal{A}}$ is a binary relation on $A$, $c^{\mathcal{A}}\in A$, $^{\mathcal{A}}$ is a binary relation on $B$, $G^{\mathcal{A}}:B\to A$, and either $\neg R(c^{\mathcal{A}})$ holds or $E^{\mathcal{A}}$ fails to satisfy the Extensionality Axiom or $^{\mathcal{A}}$ is a linear order (on $B$) without least element and $G^{\mathcal{A}}$ is order preserving]. It is elementary in $\mathcal{L}(Q_{1},\mathrm{W})$ (even $\mathcal{L}(Q_{1})$ ), and its projection yields exactly the complement model class of ${Q}_{R}$. ∎
To conclude, symbiosis is not enough to imply that the $\Delta$ operation preserves the $\mathcal{L}$-constructible universe:
Theorem 23.
There is a logic $\mathcal{L}$ which is symbiotic with a set theoretic predicate, such that $Con(\mathrm{ZF})$ implies $Con(C(\mathcal{L})\neq C(\Delta(\mathcal{L})))$.
Proof.
Let $\mathcal{L}^{*}=\mathcal{L}(Q_{1},\mathrm{W})$. Note that $\Delta(\mathcal{L}(Q_{1}))\subset\Delta(\mathcal{L}^{*})$. So, as in the proof of Theorem 16, we can force over $L$ to obtain a Cohen real $a$ and a forcing extension of $L$, say $W^{\prime}$, such that $a\in C(\Delta(\mathcal{L}^{*}))^{W^{\prime}}$, while still $C(\mathcal{L}^{*})^{W^{\prime}}=L$. ∎
5 $0^{\sharp}$ and the $\Delta$-operation
Large cardinals give an even stronger result than the above Theorem 23. Let us denote $\Delta(\mathcal{L}(Q_{1},\mathrm{W}))$ by $\Delta{(Q_{1},\mathrm{W})}$ (and similarly for $\Sigma,\Pi$) and $C(\Delta(\mathcal{L}(Q_{1},\mathrm{W})))$ by $C(\Delta(Q_{1},\mathrm{W}))$. We are indebted to Menachem Magidor for suggesting the following observation:
Theorem 24.
If $0^{\sharp}$ exists, then $0^{\sharp}\in C(\Delta(Q_{1},\mathrm{W}))$.
$0^{\sharp}$ can be defined in many equivalent ways, and we present the definition most suitable for our purposes. First recall the notion of indiscernibles.
Definition 25.
Let $\mathcal{M}=(M,E)$ be a model. A set $X\subseteq M$ is a set of indiscernibles for the model $(M,E)$ if for every $n\in\omega$, and every formula $\varphi\left(v_{1},\ldots,v_{n}\right)$, $\mathcal{M}\vDash\varphi\left(\alpha_{1},\ldots,\alpha_{n}\right)$ if and only if $\mathcal{M}\vDash\varphi\left(\beta_{1},\ldots,\beta_{n}\right)$ whenever $\alpha_{1}E\alpha_{2}E\ldots E\alpha_{n}$ and $\beta_{1}E\beta_{2}E\ldots E\beta_{n}$ are two $E$ increasing sequences of elements of $X$. We say that $X$ is a set of ordinal indiscernibles for $(M,E)$ if, in addition, $(M,E)\vDash\text{``}x\text{ is an ordinal''}$ for every $x\in X$.
Now, $0^{\sharp}$ is defined by the following fact.
Fact 26 (Silver [Sil71]).
If for some uncountable cardinal $\kappa$ there is a set of size $\kappa$ of ordinal indiscernibles for $(L_{\kappa},\in)$, then for every uncountable cardinal $\lambda$ there is a set of size $\lambda$ of ordinal indiscernibles for $(L_{\lambda},\in)$.
Furthermore, in such a case, there is a set of first-order formulas, denoted $0^{\sharp}$, such that if $X$ is a set of ordinal indiscernibles for $(L_{\kappa},\in)$ of size $\kappa$, and $\{\alpha_{k}\mid k<\omega\}\subseteq X$ is increasing, then
$0^{\sharp}=\{\varphi\mid(L_{\kappa},\in,\alpha_{k})_{k<\omega}\models\varphi\}.$
The crux of the second part of this fact is that if the assumption “for some uncountable cardinal $\kappa$ there is a set of size $\kappa$ of ordinal indiscernibles for $(L_{\kappa},\in)$” holds, then the set $0^{\sharp}$ does not depend on $\kappa$ or the specific set of indiscernibles $X$. Thus the assumption is commonly abbreviated as $0^{\sharp}$ exists”. In fact, $(L_{\kappa},\in)$ can be replaced with any well-founded model $(M,E)$ satisfying a large enough fragment of $\mathrm{ZFC}+V=L$. Let $_{L}$ denote the uniformly definable well-order of $L$. If $\mathcal{M}$ is a model of a large enough fragment of $\mathrm{ZFC}$, satisfying the sentence $V=L$, then $_{L}^{\mathcal{M}}$ is the interpretation of the definition of $_{L}$ in $\mathcal{M}$. Taking $\kappa=\omega_{1}$, note that $L_{\omega_{1}}$ is the unique (up to isomorphism) well-founded model of $V=L$ where $_{L}$ has order type $\omega_{1}$. So, if $0^{\sharp}$ exists then for every well-founded model $\mathcal{M}=(M,E)$ of $V=L$ such that $_{L}^{\mathcal{M}}$ has order type $\omega_{1}$, and first order formula $\varphi(x_{0},\ldots,x_{n-1})$ of set theory, the following are equivalent:
  1. (1)
    $\varphi\in 0^{\sharp}$;
  2. (2)
    There is an uncountable $X\subseteq M$ such that $X$ is a set of indiscernibles for $\mathcal{M}$, and there are $a_{0},\dots,a_{n-1}\in X$ such that $\mathcal{M}\models\varphi(a_{0},\dots,a_{n-1})$ ;
  3. (3)
    For every uncountable $X\subseteq M$ such that $X$ is a set of indiscernibles for $\mathcal{M}$ and for every $a_{0},\dots,a_{n-1}\in X$, $\mathcal{M}\models\varphi(a_{0},\dots,a_{n-1})$.
We will show that (2) and (3) can be expressed as $\Sigma{(Q_{1},\mathrm{W})}$ and $\Pi{(Q_{1},\mathrm{W})}$, respectively, so that (1) is in fact $\Delta{(Q_{1},\mathrm{W})}$.
Lemma 27.
If $0^{\sharp}$ exists then there is a formula $\Xi(u,v)\in\Delta{(Q_{1},\mathrm{W})}$ such that for any admissible set $N\supseteq\omega_{1}^{V}$ and any $\mathcal{M}\in N$,
$\varphi\in 0^{\sharp}\iff(N,\in)\models_{\Delta{(Q_{1},\mathrm{W})}}\Xi(\mathcal{M},\varphi).$
Proof.
We fix some notation, following Devlin’s [Dev17, section A.I.9], where a distinction is made between formulas of the “actual” language of set theory with signature $\{\in\}$, denoted by LST, and constructs which mimic these formulas inside set theory, i.e. as specific sets. The result is a set denoted $\mathscr{L}$, which we also treat as formulas, in a distinct signature $\{\epsilon\}$. Given a set $u$, $\mathscr{L}_{u}$ denotes formulas which include constant symbols for elements of $u$, such that the constant corresponding to $x\in u$ is denoted $\mathring{x}$. We follow Devlin’s convention where formulas of LST are denoted by capital Greek letters, and their counterparts in $\mathscr{L}$ by the corresponding lowercase letters. So for example if $\Phi(v)$ is an LST formula with free variable $v$, and $x\in u$, then there is a set $\varphi(\mathring{x})\in\mathscr{L}_{u}$ corresponding to $\Phi(x)$.
The results important to us are as follows:
Fact 28.
There are formulas of LST as follows:
  1. 1.
    $\operatorname{Fml}(w,u)$ : $w$ is a formula of $\mathscr{L}_{u}$;
  2. 2.
    $\operatorname{Fr}(z,w)$ : $z$ is the set of free variables in the formula $w$;
  3. 3.
    $\operatorname{Sat}(u,w)$ : $w$ is a sentence of $\mathscr{L}_{u}$ which is true in the structure $(u,\in)$ under the canonical interpretation (i.e. with $\in$ interpreting $\epsilon$ and $x$ interpreting $\mathring{x}$ for each $x\in u$ ).
These formulas can be modified to apply to general extensional structures $(u,e)$ (where $e\subseteq u\times u$) by taking care to replace every mention of $x\in y$ where $x,y\in u$ by $(x,y)\in e$. For example, a clause containing $(\exists x,y\in u)[(x\in y)$...] will be replaced by $(\exists x,y\in u)[(x,y)\in e$...]. So by following the same constructions, we obtain formulas $\operatorname{Fml}^{*}(u,e,w)$, $\operatorname{Fr}^{*}(u,e,z,w)$, $\operatorname{Sat}^{*}(u,e,w)$ such that:
Fact 29.
Let $\Phi\left(v_{0},\ldots,v_{n}\right)$ be a formula of LST, and $\varphi\left(v_{0},\ldots,v_{n}\right)$ its counterpart in $\mathscr{L}$. Then for every well-founded extensional model $\mathcal{M}=(M,E)$ and $x_{0},\dots,x_{n}\in M$,
$\mathcal{M}\models\Phi(x_{0},\dots,x_{n})\iff\operatorname{Sat}^{*}(M,E,\varphi(\mathring{x}_{0},\dots,\mathring{x}_{n}))$
Using this machinery, we will define a formula of LST capturing the notion “$X$ is a set of indiscernibles for $\mathcal{M}=(M,E)$” (assuming $\mathcal{M}$ is a well-founded extensional model). We first describe some auxiliary notation. These notions should formally refer to a pair $(M,E)$ via variables $u,e$ as before, but for the sake of readability we will not always be explicit about it. So for example, when we talk about ordinals, we mean “ordinals in the sense of $\mathcal{M}=(M,E)$”. We will also write e.g. $\psi\in\operatorname{Fml}^{*}(M,E)$ instead of $\operatorname{Fml}^{*}(M,E,\psi)$.
  • An assignment (in $\mathcal{M}$) is a function whose domain is a finite subset of variables (in the sense of $\mathcal{M}$). We use $\sigma\in\operatorname{Asn}^{*}(X,Y)$ to denote that $\sigma$ is assignments with domain exactly $X$ and range contained in $Y$.  [^2] This is in fact the same as $\sigma:X\to Y$, which we avoid to prevent confusion with the logical implication symbol $\to$ in the formulas presented below.
  • Given $\psi\in\operatorname{Fml}^{*}(M,E)$ and an assignment $\sigma$ such that the domain $\operatorname{Dom}(\sigma)$ of $\sigma$ is $\operatorname{Fr}^{*}(M,E,\psi)$, we denote by $\psi[\sigma]$ the formula resulting from $\psi$ by substituting the free variables by the constants corresponding to the values of $\sigma$. Note that this is an effective transformation on formulas in $\mathscr{L}_{u}$.
  • $v\in{\bf Ord}^{*}(M,E)$ is the formula expressing that $v$ is an ordinal in the sense of $(M,E)$.
  • If $\sigma_{1},\sigma_{2}$ are assignments with ordinal values (i.e. functions whose domain is a finite subset of variables to the ordinals of $\mathcal{M}$), we denote by $\sigma_{1}\sim\sigma_{2}$ the statement:
    $\operatorname{Dom}(\sigma_{1})=\operatorname{Dom}(\sigma_{2})\land\forall v,u\in\operatorname{Dom}(\sigma_{1})[\sigma_{1}(v)<\sigma_{1}(u)\leftrightarrow\sigma_{2}(v)<\sigma_{2}(u)].$
Now our desired formula, with predicate $X$ and parameter $\mathcal{M}$, denoted $\operatorname{Ind}(X,\mathcal{M})$, is:
$\forall v[X(v)\to v\in{\bf Ord}^{*}(M,E)]\land\forall\psi\in\operatorname{Fml}^{*}(M,E)\forall\sigma_{1},\sigma_{2}\in\operatorname{Asn}(\operatorname{Fr}^{*}(M,E,\psi),X)$
$[\sigma_{1}\sim\sigma_{2}\to(\operatorname{Sat}^{*}(M,E,\psi[\sigma_{1}])\leftrightarrow\operatorname{Sat}^{*}(M,E,\psi[\sigma_{2}]))].$
Let $\Gamma$ be the LST sentence $V=L+\mathrm{ZFC}^{-}_{n}$” (for a large enough $n$), with $\gamma$ being the corresponding $\mathscr{L}$-sentence. Let $\Upsilon(u,e,x,y)$ be the LST formula such that whenever $(M,E)$ is a model of “V=LV=L”, $\Upsilon(M,E,x,y)$ defines its canonical well-order. Given a formula $\Psi=\Psi(x,y)$ which defines a linear order, let
$\Omega xy\Psi(x,y)=Q_{1}y\exists x\Psi(x,y)\land\forall y\neg Q_{1}x\Psi(x,y)$
be the $\mathcal{L}(Q_{1})$ formula stating that the order defined by $\Psi$ is $\omega_{1}$-like.
Using this we can translate (2) and (3) above (together with the needed assumptions on $\mathcal{M}$) to the following formulas, using $\mathcal{M},\varphi$ as first-order parameters:
$2(\mathcal{M},\varphi)\quad$ $\mathcal{M}=(M,E)\land\mathrm{W}uvE(u,v)\land\operatorname{Sat}^{*}(M,E,\gamma)\land\Omega xy\Upsilon(M,E,x,y)$
$\to\ \exists X\subseteq M\big\{Q_{1}vX(v)\land\operatorname{Ind}(X,\mathcal{M})$
$\quad\land\exists\sigma\in\operatorname{Asn}(\operatorname{Fr}^{*}(M,E,\varphi),X)[\operatorname{Sat}^{*}(M,E,\varphi[\sigma])]\big\}$
$3(\mathcal{M},\varphi)\quad$ $\mathcal{M}=(M,E)\land\mathrm{W}uvE(u,v)\land\operatorname{Sat}^{*}(M,E,\gamma)\land\Omega xy\Upsilon(M,E,x,y)$
$\to\forall X\subseteq M\big\{Q_{1}vX(v)\land\operatorname{Ind}(X,\mathcal{M})$
$\quad\to\forall\sigma\in\operatorname{Asn}(\operatorname{Fr}^{*}(M,E,\varphi),X)[\operatorname{Sat}^{*}(M,E,\varphi[\sigma])]\big\}$
So, replacing $\mathcal{M},\varphi$ with variables $u,v$ (and making all the implicit definitions explicit), the above formulas are equivalent for every $u,v$. The first formula is $\Sigma(Q_{1},\mathrm{W})$ and the second $\Pi(Q_{1},\mathrm{W})$, hence they are both $\Delta(Q_{1},\mathrm{W})$. Hence, if we denote by $\Xi(u,v)$ the $\Delta{(Q_{1},\mathrm{W})}$ formula represented by this pair, we get that indeed for any admissible $N\supseteq\omega_{1}^{V}$ and any $\mathcal{M}\in N$,
$\varphi\in 0^{\sharp}\iff(N,\in)\models_{\Delta{(Q_{1},\mathrm{W})}}\Xi(\mathcal{M},\varphi).\qed$
Proof of theorem 24.
Assume $0^{\sharp}$ exists and let $C(\Delta(Q_{1},\mathrm{W}))=\bigcup L^{\prime}_{\alpha}$ be the levels of the construction, as in definition 6. Let $\alpha$ be such that $L_{\omega_{1}}\in L^{\prime}_{\alpha}$ (here $\omega_{1}=\omega_{1}^{V}$). For every first order formula $\varphi(v_{0},\dots,v_{n})$ in the signature $\{\epsilon\}$, and $x_{0},\dots,x_{n}\in L_{\omega_{1}}$ we have
$\operatorname{Sat}^{*}(L_{\omega_{1}},\in\restriction_{L_{\omega_{1}}},\varphi(\mathring{x}_{0},\dots,\mathring{x}_{n}))\iff$
$(L_{\omega_{1}},\in\restriction_{L_{\omega_{1}}})\models\Phi(x_{0},\dots,x_{n})\iff$
$(L^{\prime}_{\alpha},\in)\models\operatorname{Sat}^{*}(L_{\omega_{1}},\in\restriction_{L_{\omega_{1}}},\varphi(\mathring{x}_{0},\dots,\mathring{x}_{n})).$
For short, we denote the model $(L_{\omega_{1}},\in\restriction_{L_{\omega_{1}}})$ simply by $L_{\omega_{1}}$. So for $X\subseteq L^{\prime}_{\alpha}$, we will get that
$\operatorname{Ind}(L_{\omega_{1}},X)\iff(L^{\prime}_{\alpha},\in,X)\models\operatorname{Ind}(L_{\omega_{1}},X).$
Note that here we are considering first-order satisfaction. So now, if we consider satisfaction in $\mathcal{L}{(Q_{1},\mathrm{W})}$ and $\Delta{(Q_{1},\mathrm{W})}$, we get that for every $\varphi$
$2(L_{\omega_{1}},\varphi)$ $\iff(L^{\prime}_{\alpha},\in)\models_{\mathcal{L}{(Q_{1},\mathrm{W})}}2(L_{\omega_{1}},\varphi)$
$3(L_{\omega_{1}},\varphi)$ $\iff(L^{\prime}_{\alpha},\in)\models_{\mathcal{L}{(Q_{1},\mathrm{W})}}3(L_{\omega_{1}},\varphi),$
and by the result of the previous lemma, we get
$(L^{\prime}_{\alpha},\in)\models_{\Delta{(Q_{1},\mathrm{W})}}\Xi(L_{\omega_{1}},\varphi)\iff\varphi\in 0^{\sharp}.$
Hence
$0^{\sharp}=\{\varphi\mid(L^{\prime}_{\alpha},\in)\models_{\Delta{(Q_{1},\mathrm{W})}}\Xi(L_{\omega_{1}},\varphi)\}\in C(\Delta(Q_{1},\mathrm{W})).\qed$
6 Conclusion and open questions
Let us summarize what has been obtained:
Theorem.
  1. 1.
    $\mathrm{ZF}\vdash C(\mathcal{L}(Q_{0}))=C(\Delta(\mathcal{L}(Q_{0})))=L$ (while $\mathcal{L}(Q_{0})\neq\Delta(\mathcal{L}(Q_{0}))$).
  2. 2.
    $\mathrm{ZFC}\vdash C(\mathcal{L}^{2})=C(\Delta(\mathcal{L}^{2}))=\operatorname{HOD}$ (while $\mathcal{L}^{2}\neq\Delta(\mathcal{L}^{2})$).
  3. 3.
    It is consistent with $\mathrm{ZFC}$ that $C(\mathcal{L}(H))=\operatorname{HOD}_{1}\neq\operatorname{HOD}=C(\Delta(\mathcal{L}(H)))$.
  4. 4.
    It is consistent with $\mathrm{ZFC}$ that $C(\mathcal{L}(Q_{1}))=L\neq C(\Delta(\mathcal{L}(Q_{1})))$ .
  5. 5.
    It is consistent with $\mathrm{ZFC}$ that $C(\mathcal{L}(Q_{1},\mathrm{W}))=L\neq C(\Delta(\mathcal{L}(Q_{1},\mathrm{W})))$, where $\mathcal{L}(Q_{1},\mathrm{W})$ is symbiotic with the predicate $R_{\aleph_{1}}$.
  6. 6.
    If $0^{\sharp}$ exists, then $C(\mathcal{L}(Q_{1},\mathrm{W}))=L\neq C(\Delta(\mathcal{L}(Q_{1},\mathrm{W})))\ni 0^{\sharp}$.
This leaves the following question open:
Question 30.
Are there necessary and/or sufficient conditions on the logic $\mathcal{L}$ (besides the trivial $\mathcal{L}=\Delta(\mathcal{L})$) which imply $\mathrm{ZF}\vdash C(\mathcal{L})=C(\Delta(\mathcal{L}))$?
Dual to this would be trying to identify properties of a logic that would allow $C(\mathcal{L})\neq C(\Delta(\mathcal{L}))$:
Question 31.
Are there necessary and/or sufficient conditions on the logic $\mathcal{L}$ which imply the consistency of $C(\mathcal{L})\neq C(\Delta(\mathcal{L}))$?
The path to answering both these questions would benefit from analysing additional examples of logics and their $\Delta$-extensions. Such examples may include the logics obtained from Magidor-Malitz or cofinality quantifiers ($\mathcal{L}(Q^{\mathrm{MM}})$ and $\mathcal{L}(Q^{\mathrm{cf}}_{\alpha})$), stationary logic $\mathcal{L}(\mathtt{aa})$, and so on.
A further question regarding the theory of models constructed from logics obtained by the $\Delta$-operation is the status of AC in them. As we mentioned in the beginning (page 3), the known methods of proving that AC holds in models of the form $C(\mathcal{L})$ requires $\mathcal{L}$ to have recursive syntax, which $\Delta(\mathcal{L})$ is not known to have in general. In particular, we do not know whether the models $C(\Delta(Q_{1}))$ and $C(\Delta(Q_{1},\mathrm{W}))$ provably satisfy AC, and even whether AC holds in the specific constructions we used here. Answering these questions will require a deeper analysis of the ways forcing affects truth in the logics $\Delta{(Q_{1})}$ and $\Delta{(Q_{1},\mathrm{W})}$, which is beyond the scope of this paper. Hence the following is open:
Question 32.
What is the status of the Axiom of Choice in models of the form $C(\Delta(\mathcal{L}))$?
Another direction for further research is the impact of large cardinals on $C(\Delta(\mathcal{L}))$. Theorem 24 gives an initial result showing that large cardinals do affect the expressive power of logics obtained by the $\Delta$-operation. Hence we may ask a somewhat broad question:
Question 33.
What is the effect of large cardinals on the relationship between $C(\mathcal{L})$ and $C(\Delta(\mathcal{L}))$?
On a more concrete level, the models $C(\Delta(Q_{1}))$ and $C(\Delta(Q_{1},\mathrm{W}))$ require further investigation. For example, a first step would be to answer the following:
Question 34.
  1. 1.
    Is it consistent that $C(\Delta(Q_{1}))\neq C(\Delta(Q_{1},\mathrm{W}))$?
  2. 2.
    Do the models $C(\Delta(Q_{1})),C(\Delta(Q_{1},\mathrm{W}))$ satisfy $AC$?
  3. 3.
    Do the models $C(\Delta(Q_{1})),C(\Delta(Q_{1},\mathrm{W}))$ provably contain all the reals?
Another set of questions regards the question of idempotence: for a logic $\mathcal{L}$, we say that $C(\mathcal{L})$ is provably idempotent if it is provable that $C(\mathcal{L})^{C(\mathcal{L})}=C(\mathcal{L})$; and consistently idempotent (and $\neq L$) if it is consistent that $C(\mathcal{L})^{C(\mathcal{L})}=C(\mathcal{L})\neq L$. We will usually omit the phrase in parentheses. Note that the last statement is equivalent to “it is consistent that $V=C(\mathcal{L})\neq L$ holds”. Of course it is only interesting in the case that consistently, $C(\mathcal{L})\neq L$, as $L$ is provably idempotent. Other examples of provable idempotence are models constructed from infinitary logics $C(\mathcal{L}_{\kappa,\kappa})$ [Cha71], but this is the exception rather than the rule, as models such as $\operatorname{HOD},C^{*},C(\mathtt{aa})$ are only consistently idempotent (see [McA71, McA74, KMV21, KMV25, Ya’23, Ya’25]).
It is likely that the constructions we produced here for the logics $C(\Delta(Q_{1}))$ and $C(\Delta(Q_{1},\mathrm{W}))$ result in models where idempotence fails, as, intuitively, we only code the truth of some fact (for example which trees are not Aronszajn), but not the witnesses to this fact (the generic branches). However proving that this is in fact the case requires further research, and we leave the following open:
Question 35.
Are the models $C(\Delta(Q_{1}))$ and $C(\Delta(Q_{1},\mathrm{W}))$ consistently or even provably idempotent?
Many more questions may be asked, which we hope to address in further research.
6.1 The $\Delta_{1}^{1}$ operation
Finally, we mention a notion closely related to the $\Delta$-operation – the $\Delta_{1}^{1}$-operation – defined in the same way, except that when expanding the vocabulary we only allow adding new predicates, not new sorts. This results in a weaker closure operation on logics, as for example $\Delta_{1}^{1}(\mathcal{L}^{2})=\mathcal{L}^{2}\neq\Delta(\mathcal{L}^{2})$ (for more on this, see [Vää83]). Note that in this case,
$\operatorname{HOD}=C(\mathcal{L}^{2})=C(\Delta_{1}^{1}(\mathcal{L}^{2}))=C(\Delta(\mathcal{L}^{2})).$
However the equality $C(\Delta_{1}^{1}(\mathcal{L}))=C(\Delta(\mathcal{L}))$ is not true in general, as can be seen in the following example. Recall Theorem 15 where we show that consistently $C(\mathcal{L}(H))\neq C(\mathcal{L}(\Delta(H)))$.
Proposition 36.
$C(\Delta_{1}^{1}(\mathcal{L}(H)))=C(\mathcal{L}(H))$
Proof.
The direction $\supseteq$ is clear. Now note that a $\Sigma_{1}^{1}(L(H))$-formula can be expressed in the form $\exists R_{1}\dots\exists R_{n}\Phi$ with $\Phi\in L(H)$, and as we noted in the proof of Theorem 14, in the presence of a pairing function we can express the existential quantifiers in the terms of the Henkin quantifier. So we can inductively prove the other direction. ∎
Corollary 37.
It is consistent with $\mathrm{ZFC}$ that $C(\Delta_{1}^{1}(\mathcal{L}(H)))\neq C(\Delta(\mathcal{L}(H)))$.
Nevertheless, the results in sections 3-5 apply to the $\Delta_{1}^{1}$-operation as well, as the formulas used to separate the models only require adding predicates, not new sorts. Thus the following holds:
Theorem 38.
  1. 1.
    It is consistent with $\mathrm{ZFC}$ that $C(\mathcal{L}(Q_{1}))=L\neq C(\Delta_{1}^{1}(\mathcal{L}(Q_{1})))$.
  2. 2.
    It is consistent with $\mathrm{ZFC}$ that $C(\mathcal{L}(Q_{1},\mathrm{W}))=L\neq C(\Delta_{1}^{1}(\mathcal{L}(Q_{1},\mathrm{W})))$, where $\mathcal{L}(Q_{1},\mathrm{W})$ is symbiotic with the predicate $R_{\aleph_{1}}$.
  3. 3.
    If $0^{\sharp}$ exists, then $C(\mathcal{L}(Q_{1},\mathrm{W}))=L\neq C(\Delta_{1}^{1}(\mathcal{L}(Q_{1},\mathrm{W})))\ni 0^{\sharp}$.